# Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction

### v3

@@ A.8 $O(3)$ equivariance for crystal tensor properties
Due to the difference between the input space $\mathcal{V}$ and the output space $\mathcal{W}$, the $O(3)$ equivariance required for crystal tensor property prediction tasks differs from the $O(3)$ equivariance typically encountered in general molecular studies [13, 48, 40]. In molecular studies, the input space $\mathcal{V}$ and the output space $\mathcal{W}$ are the same, meaning group representations in these spaces are identical, i.e. $\rho_{\mathcal{V}}(g)=\rho_{\mathcal{W}}(g)$. Specifically, for a function $f(\mathbf{X})$, if it is $O(3)$ equivariant, then it satisfies $f(\mathbf{Q}\mathbf{X})=\mathbf{Q}f(\mathbf{X})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an arbitrary orthogonal matrix, and $\mathbf{X}\in\mathbb{R}^{3\times N}$ represents the coordinate matrix of atoms in a molecule. In this context, the function $f$ can be seen as the model. However, in the prediction of crystal tensor properties, the requirements for $O(3)$ equivariance differ. In what follows, we will provide a detailed introduction to this difference.
@@
Dielectric tensor. For the dielectric tensor $\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$ and a function $f:(\mathbf{A},\mathbf{F},\mathbf{L})\to\bm{\varepsilon}$, if it is $O(3)$ equivariant, it must satisfy $f(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top}$. The reason for this difference lies in the physical nature of the dielectric tensor, which characterizes a material’s polarization response to an external electric field [51]. Specifically, the dielectric tensor describes the relationship between the electric displacement $\mathbf{D}\in\mathbb{R}^{3}$ and the applied electric field $\mathbf{E}\in\mathbb{R}^{3}$ through the equation $\mathbf{D}=\bm{\varepsilon}\mathbf{E}$. When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal structure, the relationship $\mathbf{D^{\prime}}=\bm{\varepsilon}^{\prime}\mathbf{E^{\prime}}$ holds, where $\mathbf{D^{\prime}}=\mathbf{Q}\mathbf{D}$ and $\mathbf{E^{\prime}}=\mathbf{Q}\mathbf{E}$. Substituting these into the equation, we have:
$\mathbf{Q}\mathbf{D}=\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E}\to\mathbf{D}=\mathbf{Q}^{\top}\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E},$ (21)
@@
which implies that the dielectric tensor transforms under the $O(3)$ group transformation as: $\bm{\varepsilon}^{\prime}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}$. This transformation principle extends similarly to other crystal tensors, as follows.
@@
Piezoelectric tensor. The piezoelectric tensor $\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ to the electric displacement field $\mathbf{D}\in\mathbb{R}^{3}$ within the material. Mathematically, this relationship is expressed as $\mathbf{D}_{i}=\sum_{jk}\mathbf{e}_{ijk}\bm{\epsilon}_{jk}$, with $i,j,k\in\{1,2,3\}$.
@@
When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and electric displacement field is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\mathbf{D}_{i}^{\prime}=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$. The relation is then reformulated as $\mathbf{D}_{i}^{\prime}=\sum_{jk}\mathbf{e}_{ijk}^{\prime}\bm{\epsilon}_{jk}^{\prime}$. Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$. Consequently, $\mathbf{D}_{i}^{\prime}$ can be represented as
@@
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}\bm{\epsilon}_{jk}$
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}(\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn})$
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{mn}\mathbf{e}_{\ell mn}(\sum_{jk}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}^{\prime}_{jk})\quad(exchange\,sign,m\leftrightarrow j,n\leftrightarrow k)$
$=\sum_{jk}\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}\bm{\epsilon}_{jk}^{\prime}$
@@
Elastic tensor. The elastic tensor $C\in\mathbb{R}^{3\times 3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ and the stress tensor $\sigma\in\mathbb{R}^{3\times 3}$ within the material. This relationship is expressed as $\bm{\sigma}_{ij}=\sum_{k\ell}C_{ijk\ell}\bm{\epsilon}_{k\ell}$, with $i,j,k,\ell\in\{1,2,3,4\}$.
When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and stress tensor is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\bm{\sigma}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\sigma}_{mn}$. Under this transformation, the relationship becomes $\bm{\sigma}_{ij}^{\prime}=\sum_{k\ell}C_{ijk\ell}^{\prime}\bm{\epsilon}_{k\ell}^{\prime}$. Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$. Based on above equations, the transformed stress tensor $\bm{\sigma}_{ij}^{\prime}$ can be expressed as:
$\bm{\sigma}_{ij}^{\prime}$ $=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\bm{\sigma}_{mn}$ (24)
$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\bm{\epsilon}_{pq}$
$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\sum_{k\ell}\mathbf{Q}_{kp}\mathbf{Q}_{\ell q}\bm{\epsilon}^{\prime}_{k\ell}$
$=\sum_{k\ell}\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}\bm{\epsilon}^{\prime}_{k\ell}$
@@
A.9 Tensor properties symmetry
Tensor properties describe the material’s response to external physical fields (such as electric fields or mechanical stress). In materials with symmetry, this response must adhere to the symmetry requirements of the material. As demonstrated by Yan et al. [51], when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$. Therefore, the corresponding tensor properties also remain unchanged, i.e., $\bm{\varepsilon}=\mathbf{R}\bm{\varepsilon}\mathbf{R}^{\top}$. Thus, crystal symmetry imposes strict constraints on the components of the tensor, leading to the simplification or elimination of many components, reducing the number of independent components.
@@ A.9 Tensor properties symmetry
Cubic 1 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Tetragonal & Hexagonal & Trigonal 2 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Orthorhombic 3 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{22}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Monoclinic 4 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&\bm{\varepsilon}_{13}\\ 0&\bm{\varepsilon}_{22}&0\\ \bm{\varepsilon}_{13}&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Triclinic 6 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&\bm{\varepsilon}_{12}&\bm{\varepsilon}_{13}\\ \bm{\varepsilon}_{12}&\bm{\varepsilon}_{22}&\bm{\varepsilon}_{23}\\ \bm{\varepsilon}_{13}&\bm{\varepsilon}_{23}&\bm{\varepsilon}_{33}\end{pmatrix}$
@@
Independent components in the dielectric tensor. The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry. The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 8 [28].
Voigt notation for elastic tensor. Similar to the dielectric tensor, the elastic tensor also has independent elements for various types of systems due to the crystal symmetry. The elastic tensor has a minimum of 3 and a maximum of 21 independent elements for various types of systems. Voigt notation is a compact way to represent these independent components of tensor properties [14]. According to the rules $11\to 1\mathrm{~{};~{}}22\to 2\mathrm{~{};~{}}33\to 3\mathrm{~{};~{}}23,~{}32\to 4\mathrm{~{};~{}}31,~{}13\to 5\mathrm{~{};~{}}12,~{}21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix [46, 37]:
@@
The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to Wen et al. [46], Ran et al. [37]).
@@
Voigt notation for piezoelectric tensor. The number of independent components in the piezoelectric tensor for different crystal systems is shown in Tabel 10 (partial data shown; for more details, refer to De Jong et al. [4], Gorfman and Zhang [8]).
@@
A.10 Experimental details
Datasets details. The dataset for dielectric tensor and piezoelectric tensor is derived from the data processed by Yan et al. [51], sourced from the JARVIS-DFT database. Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/. The statistical details of the datasets are presented in Table 11. Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix. Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.9 for details) and then reconstruct the entire $3\times 3$ symmetric matrix. In the piezoelectric tensor dataset, the piezoelectric tensor is represented using Voigt notation as a $3\times 6$ matrix, rather than a $3\times 3\times 3$ third-order tensor (see Appendix A.9 for details). we only predict the $3\times 6$ matrix. In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor [14, 46]. Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.9 for details).

### v4

@@ A.4 $O(3)$ equivariance for crystal tensor properties
Due to the difference between the input space $\mathcal{V}$ and the output space $\mathcal{W}$, the $O(3)$ equivariance required for crystal tensor property prediction tasks differs from the $O(3)$ equivariance typically encountered in general molecular studies (hoogeboom2022equivariant; xuequivariant; song2024equivariant). In molecular studies, the input space $\mathcal{V}$ and the output space $\mathcal{W}$ are the same, meaning group representations in these spaces are identical, i.e. $\rho_{\mathcal{V}}(g)=\rho_{\mathcal{W}}(g)$. Specifically, for a function $f(\mathbf{X})$, if it is $O(3)$ equivariant, then it satisfies $f(\mathbf{Q}\mathbf{X})=\mathbf{Q}f(\mathbf{X})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an arbitrary orthogonal matrix, and $\mathbf{X}\in\mathbb{R}^{3\times N}$ represents the coordinate matrix of atoms in a molecule. In this context, the function $f$ can be seen as the model. However, in the prediction of crystal tensor properties, the requirements for $O(3)$ equivariance differ. In what follows, we will provide a detailed introduction to this difference.
@@
Dielectric tensor. For the dielectric tensor $\boldsymbol{\varepsilon}\in\mathbb{R}^{3\times 3}$ and a function $f:(\mathbf{A},\mathbf{F},\mathbf{L})\to\boldsymbol{\varepsilon}$, if it is $O(3)$ equivariant, it must satisfy $f(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top}$. The reason for this difference lies in the physical nature of the dielectric tensor, which characterizes a material’s polarization response to an external electric field (yanspace). Specifically, the dielectric tensor describes the relationship between the electric displacement $\mathbf{D}\in\mathbb{R}^{3}$ and the applied electric field $\mathbf{E}\in\mathbb{R}^{3}$ through the equation $\mathbf{D}=\boldsymbol{\varepsilon}\mathbf{E}$. When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal structure, the relationship $\mathbf{D^{\prime}}=\boldsymbol{\varepsilon}^{\prime}\mathbf{E^{\prime}}$ holds, where $\mathbf{D^{\prime}}=\mathbf{Q}\mathbf{D}$ and $\mathbf{E^{\prime}}=\mathbf{Q}\mathbf{E}$. Substituting these into the equation, we have:
$\mathbf{Q}\mathbf{D}=\boldsymbol{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E}\to\mathbf{D}=\mathbf{Q}^{\top}\boldsymbol{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E},$ (21)
@@
which implies that the dielectric tensor transforms under the $O(3)$ group transformation as: $\boldsymbol{\varepsilon}^{\prime}=\mathbf{Q}\boldsymbol{\varepsilon}\mathbf{Q}^{\top}$. This transformation principle extends similarly to other crystal tensors, as follows.
@@
Piezoelectric tensor. The piezoelectric tensor $\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$ describes the relationship between the applied strain $\boldsymbol{\epsilon}\in\mathbb{R}^{3\times 3}$ to the electric displacement field $\mathbf{D}\in\mathbb{R}^{3}$ within the material. Mathematically, this relationship is expressed as $\mathbf{D}_{i}=\sum_{jk}\mathbf{e}_{ijk}\boldsymbol{\epsilon}_{jk}$, with $i,j,k\in\{1,2,3\}$.
@@
When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and electric displacement field is transformed to $\boldsymbol{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\epsilon}_{mn}$ and $\mathbf{D}_{i}^{\prime}=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$. The relation is then reformulated as $\mathbf{D}_{i}^{\prime}=\sum_{jk}\mathbf{e}_{ijk}^{\prime}\boldsymbol{\epsilon}_{jk}^{\prime}$. Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\boldsymbol{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\boldsymbol{\epsilon}^{\prime}_{mn}$. Consequently, $\mathbf{D}_{i}^{\prime}$ can be represented as
@@
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}\boldsymbol{\epsilon}_{jk}$
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}(\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\boldsymbol{\epsilon}^{\prime}_{mn})$
$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{mn}\mathbf{e}_{\ell mn}(\sum_{jk}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\epsilon}^{\prime}_{jk})\quad(exchange\,sign,m\leftrightarrow j,n\leftrightarrow k)$
$=\sum_{jk}\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}\boldsymbol{\epsilon}_{jk}^{\prime}$
@@
Elastic tensor. The elastic tensor $C\in\mathbb{R}^{3\times 3\times 3\times 3}$ describes the relationship between the applied strain $\boldsymbol{\epsilon}\in\mathbb{R}^{3\times 3}$ and the stress tensor $\sigma\in\mathbb{R}^{3\times 3}$ within the material. This relationship is expressed as $\boldsymbol{\sigma}_{ij}=\sum_{k\ell}C_{ijk\ell}\boldsymbol{\epsilon}_{k\ell}$, with $i,j,k,\ell\in\{1,2,3,4\}$.
When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and stress tensor is transformed to $\boldsymbol{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\epsilon}_{mn}$ and $\boldsymbol{\sigma}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\sigma}_{mn}$. Under this transformation, the relationship becomes $\boldsymbol{\sigma}_{ij}^{\prime}=\sum_{k\ell}C_{ijk\ell}^{\prime}\boldsymbol{\epsilon}_{k\ell}^{\prime}$. Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\boldsymbol{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\boldsymbol{\epsilon}^{\prime}_{mn}$. Based on above equations, the transformed stress tensor $\boldsymbol{\sigma}_{ij}^{\prime}$ can be expressed as:
$\boldsymbol{\sigma}_{ij}^{\prime}$ $=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\boldsymbol{\sigma}_{mn}$ (24)
$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\boldsymbol{\epsilon}_{pq}$
$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\sum_{k\ell}\mathbf{Q}_{kp}\mathbf{Q}_{\ell q}\boldsymbol{\epsilon}^{\prime}_{k\ell}$
$=\sum_{k\ell}\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}\boldsymbol{\epsilon}^{\prime}_{k\ell}$
@@
A.5 Tensor properties symmetry
Tensor properties describe the material’s response to external physical fields (such as electric fields or mechanical stress). In materials with symmetry, this response must adhere to the symmetry requirements of the material. As demonstrated by yanspace, when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$. Therefore, the corresponding tensor properties also remain unchanged, i.e., $\boldsymbol{\varepsilon}=\mathbf{R}\boldsymbol{\varepsilon}\mathbf{R}^{\top}$. Thus, crystal symmetry imposes strict constraints on the components of the tensor, leading to the simplification or elimination of many components, reducing the number of independent components.
@@ A.5 Tensor properties symmetry
Cubic 1 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
Tetragonal & Hexagonal & Trigonal 2 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
Orthorhombic 3 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{22}&0\\ 0&0&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
Monoclinic 4 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&\boldsymbol{\varepsilon}_{13}\\ 0&\boldsymbol{\varepsilon}_{22}&0\\ \boldsymbol{\varepsilon}_{13}&0&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
Triclinic 6 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&\boldsymbol{\varepsilon}_{12}&\boldsymbol{\varepsilon}_{13}\\ \boldsymbol{\varepsilon}_{12}&\boldsymbol{\varepsilon}_{22}&\boldsymbol{\varepsilon}_{23}\\ \boldsymbol{\varepsilon}_{13}&\boldsymbol{\varepsilon}_{23}&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
@@
Independent components in the dielectric tensor. The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry. The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 7 (mao2024dielectric).
Voigt notation for elastic tensor. Similar to the dielectric tensor, the elastic tensor also has independent elements for various types of systems due to the crystal symmetry. The elastic tensor has a minimum of 3 and a maximum of 21 independent elements for various types of systems. Voigt notation is a compact way to represent these independent components of tensor properties (itin2013constitutive). According to the rules $11\to 1\mathrm{~;~}22\to 2\mathrm{~;~}33\to 3\mathrm{~;~}23,~32\to 4\mathrm{~;~}31,~13\to 5\mathrm{~;~}12,~21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix (wen2024equivariant; ran2023velas):
@@
The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 8 (partial data shown; for more details, refer to wen2024equivariant; ran2023velas).
@@
Voigt notation for piezoelectric tensor. The number of independent components in the piezoelectric tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to de2015database; gorfman2024piezoelectric).
@@
A.6 Experimental details
Datasets details. The dataset for dielectric tensor and piezoelectric tensor is derived from the data processed by yanspace, sourced from the JARVIS-DFT database. Since nearly half of the crystal samples in the piezoelectric dataset have piezoelectric tensor labels that are entirely zero, training GoeCTP directly on this dataset can lead to severe overfitting, thereby compromising the reliability of the experimental results. Therefore, we filtered out all crystal samples with zero-valued piezoelectric tensor labels from piezoelectric dataset. For the elastic tensor prediction task, the dataset is also obtained from the JARVIS-DFT database and is publicly accessible as dft_3d through the jarvis-tools package  [^3] https://pages.nist.gov/jarvis/databases/. The statistical details of the datasets are presented in Table 10. Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix. Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.5 for details) and then reconstruct the entire $3\times 3$ symmetric matrix. In the piezoelectric tensor dataset, the piezoelectric tensor is represented using Voigt notation as a $3\times 6$ matrix, rather than a $3\times 3\times 3$ third-order tensor (see Appendix A.5 for details). we only predict the $3\times 6$ matrix. In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor (itin2013constitutive; wen2024equivariant). Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.5 for details).